Find the transformation of the equation 12π‘₯2 βˆ’ 2𝑦2 + 𝑧2 = 2π‘₯𝑦 if the origin is kept fixed and the axes are rotated in such a way that the direction ratios of the new axes are 1, βˆ’3, 0; 3, 1, 0; 0, 0, 1


1
Expert's answer
2022-02-14T11:25:38-0500

Solution:

The normalized axes are



"\mathbf{e}'_1 = \frac{1}{\sqrt{10}} (1, -3, 0), \hspace{3mm} \mathbf{e}'_2 = \frac{1}{\sqrt{10}} (3, 1, 0), \hspace{3mm}
\mathbf{e}'_3 = (0, 0, 1)"

or



"\begin{pmatrix}
 \mathbf{e}'_1 \\ \mathbf{e}'_2 \\ \mathbf{e}'_3
\end{pmatrix} = 
\begin{pmatrix}
\frac{1}{\sqrt{10}} & \frac{-3}{\sqrt{10}} & 0 \\
\frac{3}{\sqrt{10}} & \frac{1}{\sqrt{10}} & 0\\
0 & 0 & 1 \\
\end{pmatrix}
\begin{pmatrix}
 \mathbf{e}_1 \\ \mathbf{e}_2 \\ \mathbf{e}_3
\end{pmatrix}"

- the basis in the rotated frame. The transformation of coordinates is given by

"\begin{pmatrix}
 x' \\ y' \\ z'
\end{pmatrix} =\begin{pmatrix}
\frac{1}{\sqrt{10}} & \frac{-3}{\sqrt{10}} & 0 \\
\frac{3}{\sqrt{10}} & \frac{1}{\sqrt{10}} & 0\\
0 & 0 & 1 \\
\end{pmatrix}^{-1}
\begin{pmatrix}
 x \\ y \\ z
\end{pmatrix}"


so



"\begin{pmatrix}
 x \\ y \\ z
\end{pmatrix} =\begin{pmatrix}
\frac{1}{\sqrt{10}} & \frac{-3}{\sqrt{10}} & 0 \\
\frac{3}{\sqrt{10}} & \frac{1}{\sqrt{10}} & 0\\
0 & 0 & 1 \\
\end{pmatrix}
\begin{pmatrix}
 x' \\ y' \\ z'
\end{pmatrix}""\bigg\{\begin{matrix}
x &=& \frac{x'}{\sqrt{10}} - \frac{3y'}{\sqrt{10}} \\
y &=& \frac{3x'}{\sqrt{10}} + \frac{y'}{\sqrt{10}} \\
z &=& z'
\end{matrix}""12x^2 -2y^2+z^2=2xy""12\frac{(x' - 3y')^2}{10} - 2\frac{(3x' +y')^2}{10} + z'^2 = 2\frac{(x' - 3y')(3x' +y')}{10}"





"12(x'^2 - 6x'y'+9y'^2) - 2(9x'^2 +6x'y' + y'^2)+ \\ + 10z'^2 = 2(3x'^2 -8x'y'-3y'^2)"




"-12x'^2 -68x'y' +112y'^2 +10z'^2 = 0"


Answer:


"5z^2 +56y'^2-6x'^2=34x'y'"